OpenStudy (anonymous):

What is the approximate value of if Theta if cos Theta=8/15?

jimthompson5910 (jim_thompson5910):

Hint: if cos(theta) = 8/15 then theta = arccos(8/15) or theta = -arccos(8/15)

OpenStudy (anonymous):

how do you put arc

OpenStudy (anonymous):

into a calculator?

jimthompson5910 (jim_thompson5910):

above the "cosine" or "cos" key is the \(\Large \cos^{-1}\) symbol, see this?

OpenStudy (anonymous):

oh so arccos means cos-1?

jimthompson5910 (jim_thompson5910):

yes, \(\Large \cos^{-1}\) means "inverse cosine"

jimthompson5910 (jim_thompson5910):

inverse cosine = arccosine

jimthompson5910 (jim_thompson5910):

so you would hit "2nd" then "cos" then type in 8/15 then close off the parenthesis group

jimthompson5910 (jim_thompson5910):

and hit enter

OpenStudy (anonymous):

i got 1.00

OpenStudy (anonymous):

but its not a answer choice?

jimthompson5910 (jim_thompson5910):

make sure you're in degree mode

jimthompson5910 (jim_thompson5910):

do you know how to change from degrees to radians and vice versa?

OpenStudy (anonymous):

i got 57.76

jimthompson5910 (jim_thompson5910):

me too so that's one value of theta

jimthompson5910 (jim_thompson5910):

another value of theta is -57.76

OpenStudy (anonymous):

one of the answer choices is 58 so this would be the answer correct?

OpenStudy (anonymous):

bc it says approximately

jimthompson5910 (jim_thompson5910):

yeah that's close enough what do the other ones say?

OpenStudy (anonymous):

32 28 62 48

jimthompson5910 (jim_thompson5910):

ok so it's 58 then

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

there are infinitely many of such values, but none of them are listed (except for 58 of course)

jimthompson5910 (jim_thompson5910):

np